OpenStudy (anonymous):

Need Help asap!!!!! Select the ordered pair from the choices below that is a solution to the following system of equations: 3x = 2y − 16 4x − y = -13 Need Help!!!!!!

OpenStudy (allank):

There are three methods you can use here: substitution, elimination and augmented matrix. Which do you want help with?

OpenStudy (ghazi):

multiply second equation by 2 and add it to the first you'll get x and hence y

OpenStudy (allank):

@ghazi has outlined elimination. For substitution, you can make y the subject of the second equation. y=4x+13 then plug that into the first equation to get the value of y.

OpenStudy (anonymous):

(-4, 2) (-5, 1/2) (8, 12) (-2, 5) its one of these :/

OpenStudy (ghazi):

solve it first

OpenStudy (allank):

Before we tell you the answer, work it out and tell us yours.

OpenStudy (anonymous):

-_-

OpenStudy (ghazi):

\[3x-2y=-16, (4x-y=-13)*2= 8x-2y=-26\]

OpenStudy (ghazi):

now subtract 3x-2y=-16 from 8x-2y=-26 , you'll get x= -2 and y=5